# Motion Focus Recognition in Fast-Moving Egocentric Video

### v1

@@ Abstract
From Vision–Language–Action (VLA) systems to robotics, existing egocentric datasets primarily focus on action recognition tasks, while largely overlooking the inherent role of motion analysis in sports and other fast-movement scenarios. To bridge this gap, we propose a real-time motion focus recognition method that estimates the subject’s locomotion intention from any egocentric video. Our approach leverages the foundation model for camera pose estimation and introduces system-level optimizations to enable efficient and scalable inference. Evaluated on a collected egocentric action dataset, our method achieves real-time performance with manageable memory consumption through a sliding batch inference strategy. This work makes motion-centric analysis practical for edge deployment and offers a complementary perspective to existing egocentric studies on sports and fast-movement activities. Our code and dataset are available on: https://arazi2.github.io/aisends.github.io/project/VisionGPT
@@ 3.2 Motion Focus Recognition
To enable motion-aware prioritization in fast movements, we introduce a lightweight Motion Focus Recognition module that estimates the impact region induced by the user’s locomotion. The core intuition is that the acceleration vector induced by camera translation, when projected onto the image plane, provides a physically grounded focus of attention indicating where near-field interactions or potential collisions are most likely to occur [7, 1, 8].
@@
Figure 4: Motion focus calculation process. We first calculate the camera center and project the acceleration vector to the $k$th frame. Then we aggregate $K$ camera centers and apply Gaussian kernels for rendering. The final results represent the camera movement trends, which aligned with the defined motion focus.
@@
As shown in Figure 4, we then construct a motion focus map by aggregating $K$ (e.g., 15) motion points in the image plane. Each motion point contributes a Gaussian kernel centered at its projected pixel location, with the kernel spread modulated by the corresponding motion magnitude to reflect its spatial influence. The final focus map is obtained by summing all kernels and normalizing the result:
$M(u,v)=\frac{1}{Z}\sum_{i=1}^{K}\exp\!\left(-\frac{(u-u_{i})^{2}+(v-v_{i})^{2}}{2(\sigma s_{i})^{2}}\right),$ (7)
@@
The resulting motion focus is characterized by both its spatial location in the image and its associated depth and forward velocity, allowing the system to prioritize recognition in regions that align with the predicted future path.

### v2

@@ Abstract
From Vision–Language–Action (VLA) systems to robotics, existing egocentric datasets primarily focus on action recognition tasks, while largely overlooking the inherent role of motion analysis in sports and other fast-movement scenarios. To bridge this gap, we propose a real-time motion focus recognition method that estimates the subject’s locomotion intention from any egocentric video. We leverages the foundation model for camera pose estimation and introduces system-level optimizations to enable efficient and scalable inference. Evaluated on a collected egocentric action dataset, our method achieves real-time performance with manageable memory consumption through a sliding batch inference strategy. This work makes motion-centric analysis practical for edge deployment and offers a complementary perspective to existing egocentric studies on sports and fast-movement activities. Our code and dataset are available on: https://arazi2.github.io/aisends.github.io/project/VisionGPT
@@ 2.3 Motion Focus Recognition
To enable motion-aware prioritization in fast movements, we introduce a lightweight Motion Focus Recognition module that estimates the impact region induced by the user’s locomotion. The core intuition is that the acceleration vector induced by camera translation, when projected onto the image plane, provides a physically grounded focus of attention indicating where near-field interactions or potential collisions are most likely to occur [7, 1].
@@
Figure 4: Motion focus calculation process. We first calculate the camera center and project the acceleration vector to the $n$-th frame. Then we aggregate NN camera centers and apply Gaussian kernels for rendering. The final results represent the camera movement trends, which aligned with the defined motion focus.
@@
As shown in Figure 4, we then construct a motion focus map by aggregating $N$ motion points (e.g., $N=15$) in the image plane. Each motion point contributes a Gaussian kernel centered at its projected pixel location, with the kernel spread modulated by the corresponding motion magnitude to reflect its spatial influence. The final focus map is obtained by summing all kernels and normalizing the result:
$M(u,v)=\frac{1}{Z}\sum_{i=1}^{N}\exp\!\left(-\frac{(u-u_{i})^{2}+(v-v_{i})^{2}}{2(\sigma s_{i})^{2}}\right),$ (7)
@@
The resulting motion focus allowing the system to prioritize recognition in regions that align with the predicted future path.